OpenStudy (anonymous):

Derivative of f(x)= (tanx-5)/secx can someone show me a step by step

OpenStudy (amistre64):

what have you tried step by step so far?

OpenStudy (amistre64):

i assume your confusing yourself with the quotient rule

OpenStudy (amistre64):

product rule is always much easier to work with

OpenStudy (amistre64):

f(x)= (tanx-5)/secx = (tanx-5) * (secx)^-1 = (tanx-5) * (cosx) and derive

OpenStudy (amistre64):

\[D[r_1r_2r_3...r_n]=r_1'r_2r_3...r_n+r_1r_2'r_3...r_n+...+r_1r_2r_3...r_n'\]

OpenStudy (anonymous):

I'm just confused with the alegbra/ trig after using the quotient rule

OpenStudy (anonymous):

I dont see how you are suppose to get secx-sinx(-5+tanx)

OpenStudy (amistre64):

(tanx-5) * (cosx) => sec^2(x) cos(x) + (tan(x)-5)(-sin(x)) => sec(x) - tan(x)sin(x) -5sin(x)

OpenStudy (amistre64):

dont use the quotient rule; it just adds to much extra confusion that is not needed

OpenStudy (anonymous):

makes way more sense >> lol